# Thermodynamics Reconstructed from Information Science: Theoretical Core

### v1

@@ 6 Relation to existing theories and features of the present framework
This section summarizes how our theoretical core relates to existing information thermodynamics and highlights conceptual features and open problems.
6.1 Relation to existing information thermodynamics
Expressing dissipation as a path-space KL divergence,
@@ 6.2 Conceptual features
    Primary definition of pressure via information-volume constraints. We introduce volume not as geometric space volume but via the cell volume $v(y)=\tau(C_{y})$ determined by $(M,\tau)$ and the associated log-volume $\ell_{V}(x)$. Incorporating the constraint $\mathbb{E}[\ell_{V}]=\bar{v}$ into Min-KL inference defines pressure $P$ as the shadow price (dual variable), on equal footing with temperature $T$ and chemical potential $\mu$ (Eq. (10)). This repositions pressure as the variable conjugate to an observation- and convention-dependent information volume, enabling unified extensions to nonstandard state spaces and coarse-grainings.
@@
    Heat reconstructed as an information-theoretic representation theorem. With dissipation defined by path KL, system entropy defined by observational entropy, and temperature defined as a Min-KL dual variable, we define $\Delta S_{\mathrm{env}}=\Sigma_{0,T}-\Delta S_{\mathrm{sys}}$ and represent heat as
    $Q=T\bigl(\Sigma_{0,T}-\Delta S_{\mathrm{sys}}\bigr).$
@@ 6.3 Open problems and directions
We conclude by listing several theoretical and applied directions.
@@
    Phase-transition structure with information-volume constraints. In Min-KL inference with $\mathbb{E}[\ell_{V}]=\bar{v}$, it is important to analyze where the dual map $(\beta,\alpha,\pi)\mapsto(U,\bar{N},\bar{v})$ loses regularity and when ensemble nonequivalence occurs, which directly relates to phase transitions involving information volume. In particular, effective models without geometric volume or with nonlocal measurements $M$ may display novel critical phenomena.
@@
    Optimizing observational dissipation bounds and observation design. The gap between observational dissipation $\Sigma_{\mathrm{obs}}$ and true dissipation $\Sigma_{0,T}$ (and hence hidden dissipation) is sensitive to $(M,\tau)$. Designing $(M,\tau)$ to maximize $\Sigma_{\mathrm{obs}}$ under experimental constraints is a natural optimization problem linking statistical efficiency of dissipation estimation to experimental design.
@@
    New decomposition formulas combining the first-law-type relation and path dissipation. Combining the static first-law-type relation (15) with path-KL dissipation $\Sigma_{0,T}$ suggests new decompositions/inequalities that simultaneously involve changes in internal energy, information volume, and particle number together with dissipation, correlation generation, and hidden dissipation. For instance, deriving second-law-type inequalities including an information-volume work-like term and the production of correlation free energy $TI(X;Y)$ under a general prior $\tau$ would further tighten the link between nonequilibrium thermodynamics and information theory.
@@
    Applications to concrete models and comparisons with existing principles. In planned applications to RNAP–DNA systems and chemical reaction systems analyzed via large deviation theory [17], explicitly evaluating correlation free energy, hidden dissipation, and the heat representation (39) is essential for comparing with fluctuation theorems and Landauer-type principles. Clarifying when conventional heat/work identifications agree with (or differ from) our information-volume/dissipation-based identifications is also crucial for interpretation in applications.
In summary, our information-volume constraint approach encompasses existing results in information thermodynamics [9, 10, 11, 12, 13, 3, 4, 5, 6, 7, 14, 15, 16, 17] while explicitly elevating $(M,\tau)$ to primitive building blocks. This repositions static and dynamic information thermodynamics within a single KL-geometric framework and provides an abstract foundation for future applications and experimental connections.

### v2

@@ 6 Relation to existing theories and features of the present framework
This section summarizes how our theoretical core relates to existing information thermodynamics and highlights conceptual features and open problems. Expressing dissipation as a path-space KL divergence,
@@ 6.1 Conceptual features
    Primary definition of pressure via information-volume constraints. We introduce volume not as geometric space volume but via the cell volume $v(y)=\tau(C_{y})$ determined by $(M,\tau)$ and the associated log-volume $\ell_{V}(x)$. Incorporating the constraint $\mathbb{E}[\ell_{V}]=\bar{\ell}_{V}$ into Min-KL inference defines pressure $P$ as the shadow price (dual variable), on equal footing with temperature $T$ and chemical potential $\mu$ (Eq. (8)). This repositions pressure as the variable conjugate to an observation- and convention-dependent information volume, enabling unified extensions to nonstandard state spaces and coarse-grainings.
@@
    Heat reconstructed as an information-theoretic representation theorem. With dissipation defined by path KL, system entropy defined by observational entropy, and temperature defined as a Min-KL dual variable, we define
    $\Delta S_{\mathrm{env}}=\Sigma_{0,T}-\Delta S_{\mathrm{sys}},\qquad Q=T\Delta S_{\mathrm{env}}.$
@@ 6.2 Open problems and directions
Several theoretical and applied directions arise naturally.
@@
    Phase-transition structure with information-volume constraints. In Min-KL inference with $\mathbb{E}[\ell_{V}]=\bar{\ell}_{V}$, it is important to analyze where the dual map $(\beta,\alpha,\pi)\mapsto(U,\bar{N},\bar{\ell}_{V})$ loses regularity and when ensemble nonequivalence occurs, directly relating to phase transitions involving information volume. Effective models without geometric volume or with nonlocal measurements $M$ may display novel critical phenomena.
@@
    Optimizing observational dissipation bounds and observation design. The gap between observational dissipation $\Sigma_{\mathrm{obs}}$ and true dissipation $\Sigma_{0,T}$ (hence hidden dissipation) is sensitive to $(M,\tau)$. Designing $(M,\tau)$ to maximize $\Sigma_{\mathrm{obs}}$ under experimental constraints is a natural optimization problem linking statistical efficiency of dissipation estimation to experimental design.
@@
    New decomposition formulas combining the first-law-type relation and path dissipation. Combining the static first-law-type relation (13) with path-KL dissipation suggests new decompositions/inequalities involving changes in internal energy, information volume (or its code length), and particle number together with dissipation, correlation generation, and hidden dissipation. For instance, deriving second-law-like inequalities including an information-volume work-like term and the production of correlation free energy $TI(X;Y)$ under a general prior $\tau$ would deepen the link between nonequilibrium thermodynamics and information theory.
@@
    Applications to concrete models and comparison with existing principles. In planned applications to RNA polymerase–DNA systems and chemical reaction networks analyzed via large-deviation theory [17], explicit computation of correlation free energy, hidden dissipation, and heat representation (31) will clarify agreement or deviation from conventional thermodynamic descriptions such as fluctuation theorems and Landauer-type principles.
Closing remarks. Taken together, these contributions recast pressure, dissipation, and heat as quantities that can be defined and compared across descriptions through the primitive observational data $(M,\tau)$ and path-space relative entropy. By making the information-volume constraint and the path-KL definition of dissipation primary, we obtain a unified and representation-theoretic framework in which conventional thermodynamic notions emerge as secondary objects, while projection-induced gaps quantify the cost of partial observation. We expect that this viewpoint will be especially useful for systems where geometric volume, energy-bath idealizations, or microscopic reversibility assumptions are not directly available, and for designing measurements that make dissipation experimentally accessible.